OpenStudy (anonymous):

Evaluate this expression. 10 + 8 divide by 4 - 12 =

OpenStudy (anonymous):

10+8/4-12 PEMDAS 10+2-12 12-12 0

OpenStudy (anonymous):

\[\frac{ 18 }{ -8 }\]

OpenStudy (anonymous):

It depends what you're looking for. \[\frac{ 10 + 8 }{ 4 - 12 } = \frac{ 18 }{ -8 } = -2.25\]